How do I calculate price at equilibrium, quantity at equilibrium, and consumer surplus at equilibrium? Three questions essentially, please help! Thank you!
Participating in the market for a product are 12 consumers, each of whom will buy at most one unit of the good, and 11 suppliers, each of whom will sell at most one unit of the good. The distribution of buyer values is as follows: Buyer Value $1 Number of Buyers $3 4 $5 $9 3 The distribution of seller costs is as follows Seller Cost Number of Sellers $2 $4 $6 4 How many units of the good will be produced in a competitive equilibrium?
The distribution of seller costs is as follows: Seller Cost $2 $4 $6 How many units of the good will be produced in a competitive equilibrium? Number of Sellers 4 0 3 O 2
Answer #1

Equilibrium quantity=6 and equilibrium price will be 4

and consumer surplus is the difference between the willingness to pay and the actual pay

thus consumer surplus=3(5)+3(1)=15+3=18